OpenStudy (anonymous):

WILL GIVE REWARD aka medal pleaaaze helpu! 1)Simplify 3-2i/ 1-i a)1+i/2 b)1-i/2 c)5+i/2 d) no solution

OpenStudy (michele_laino):

hint: if we multiply both numerator and denominator by 1+i, we get: \[\Large \frac{{3 - 2i}}{{1 - i}} = \frac{{\left( {3 - 2i} \right)\left( {1 + i} \right)}}{{\left( {1 - i} \right)\left( {1 + i} \right)}}\] please simplify

OpenStudy (anonymous):

is it answer c?

OpenStudy (michele_laino):

yes! correct!